OpenStudy (anonymous):

How do you find the derivative of 1+2^(4-t) ?

OpenStudy (anonymous):

the one will disappear. you can set 4-t=u so you have 2^u

OpenStudy (anonymous):

then you would get ln (2)*2^u*du ln (2)*2^(4-t)*(-1) -ln (2)(2^(4-t))

OpenStudy (kira_yamato):

Let u = 4 - t du/dx = -1 dy/dx = 2^(4-t) ln 2 (-1) = -(ln2)[2^(4-t)]

OpenStudy (anonymous):

and the answer gets repeated.

OpenStudy (anonymous):